If the standard deviation of hole diameter exceeds 0.01 millimeters, there is an unacceptably high probability that the rivet will not fit. Suppose that in a sample of 15, the standard deviation is 0.008 millimeter. Is there strong evidence to indicate that the standard deviation of hole diameter exceeds 0.01 millimeter? Use α = 0.10.


1
Expert's answer
2021-09-17T03:54:19-0400

Given that "n=15" and "\sigma=0.008"

"H_o=s =0.01"

"H_1=s >0.01"


By using chi-square test:

"\frac{\left(n-1\right)\left(s^2\right)}{\left(s^2\right)}"

"=\frac{14\left(0.008\times 0.008\right)}{0.0001}"

"=8.96"

The tabulated value of chi-square deg. of freedom at 14 is 6.57.

Here calculated chisquare> tabulated chi-square.

Therefore, we reject the null hypothesis.